### v3

# Distribution Matching Distillation Meets Reinforcement Learning

Abstract
Distribution Matching Distillation (DMD) distills a pre-trained multi-step diffusion model to a few-step one to improve inference efficiency. However, the performance of the latter is often capped by the former. To circumvent this dilemma, we propose DMDR, a novel framework that combines Reinforcement Learning (RL) techniques into the distillation process. We show that for the RL of the few-step generator, the DMD loss itself is a more effective regularization compared to the traditional ones. In turn, RL can help to guide the mode coverage process in DMD more effectively. These allow us to unlock the capacity of the few-step generator by conducting distillation and RL simultaneously. Meanwhile, we design the dynamic distribution guidance and dynamic renoise sampling training strategies to improve the initial distillation process. The experiments demonstrate that DMDR can achieve leading visual quality, prompt coherence among few-step methods, and even exhibit performance that exceeds the multi-step teacher.
1 Introduction
Diffusion models have achieved unparalleled quality in visual generation tasks [dit, sd3, flux, lumina2, sdxl]. Yet their sampling process typically demands tens of iterative denoising steps, each involving a full forward pass through a large neural network, which makes high-resolution text-to-image synthesis costly both in time and computation.
To accelerate sampling speed, a surge of work [diffinstruct, dmd, lcm, hypersd, dmd2, flashdiff] has sought to distill the original diffusion model into a generator that needs only a handful of steps (e.g, 4). These approaches can be roughly divided into two types: distillation from the trajectory and distillation from the distribution. The former method operates on the instance level, aiming to simulate the original teacher diffusion generation trajectories in as few steps as possible [lcm, piflow]. While the latter works at the distribution level, striving to match the output distribution of the few-step student generator to the original multi-step one’s [diffinstruct, dmd].
In this study, we focus on distribution matching distillation (DMD) as it has been verified in large-scale settings [dmd2, dmdx] and has already been used in some leading products (e.g, Seedream 4.0 [seedream4]). However, one notable thing is that in DMD, the student model aims to match the distribution of the teacher model, which also means that the ability of the student model is often limited by the teacher. To remedy this, most works (e.g., DMD2 [dmd2]) introduce an additional GAN branch for guidance [gan]. However, training with GAN not only renders the training process dependent on external high-quality visual data, but also introduces training instability inherent in GAN [wgan, improvedgan].
To this end, we propose DMDR, which combines distribution matching distillation (DMD) with recent powerful diffusion reinforcement learning (RL) [refl, flowgrpo, diffusiondpo, ddpo]. In a nutshell, DMDR enables the student model to obtain supervision signals beyond the teacher model without relying on external image data, and makes it possible for the student to surpass the teacher. Contrary to the prevailing belief that RL should be conducted after obtaining the few-step generator, we perform RL concurrently with the distillation process. We show that such a combination is mutually beneficial: (1) As for DMD, RL can help to reshape some existing low-probability modes by guiding them toward higher reward regions, ensuring these modes are covered during the distillation process, thereby mitigating the risk of zero forcing. (2) As for RL, DMD loss itself can serve as a more effective regularization than its traditional counterparts, thereby mitigating the risk of reward hacking.
Meanwhile, since obtaining valid reward scores requires the few-step model to possess a basic generation capability already, we design two dynamic training strategies to facilitate a faster and better “cold start” for DMDR: (1) DynaDG, which strengthens the initial distributional overlap between the teacher and student model; (2) DynaRS, which enhances the student’s global modeling of the teacher’s distribution during the early training phase.
Taken together, our experimental results show that our method leads to state-of-the-art few-step generative models that can outperform their teacher, using as few as 4 sampling steps. Moreover, our method is not only compatible with different models (e.g., flow-based, denoising-based) but also with various RL algorithms (e.g., ReFL, DPO, GRPO). This ensures the long-term effectiveness of our method as the multi-step model and RL algorithm evolve.
In summary, our main contributions are as follows:
  • We propose DMDR, which shows that DMD and RL can be trained simultaneously with mutual benefits.
  • We design a dynamic cold start stage for faster and better distillation in the initial phase.
  • We validate our method on various models and RL algorithms with all of which achieve excellent performance.
Figure 1: More Images generated by Z-Image-Turbo [zimage] distilled through our DMDR.
2 Related Work
We highlight key related studies here, and leave the discussions of others in Appendix A.
Distribution Matching Distillation. Distribution Matching Distillation (DMD) [dmd] is the first work to successfully apply the principle score-based distillation [score-base-diff, prolificdreamer, dreamfusion] to large-scale text-to-image models. Intuitively, it strives to ensure that any sample realized by the student at a given noise level occurs with exactly the same probability as it would under the teacher’s distribution, thereby preserving the multi-step generative priors in the few-step model. From then on, a lot of follow-up work emerged [dmd2, dmdx, tdm, sd3.5flash, f-distill], for example, DMD2 [dmd2] employs a discriminator to align the student model distribution with a specific target distribution like GAN [gan] does; TDM [tdm] incorporates DMD loss in the sampling process of the student model for better alignment; ff–distill [f-distill] replaces the original reverse Kullback–Leibler (KL) divergence to the proposed $f$-divergence for covering different divergences with different properties. Our work also starts with DMD, but we don’t focus on how to better “imitate” the teacher. Instead, we aim to incorporate reinforcement learning in the distillation process to enable the student to outperform the teacher.
Reinforce Learning for Diffusion Models. Reinforce learning helps to align diffusion models to human preferences by training a reward model and using it to guide generation [ddpo, flowgrpo, refl, srpo, diffusiondpo, dancegrpo]. There are many algorithms to conduct RL, for example, DDPO [ddpo] adapts PPO via image log-likelihoods; ReFL [refl] bypasses likelihoods by optimizing outputs with frozen-reward gradients; Diffusion-DPO [diffusiondpo] adapts DPO to diffusion for paired human preference data; and recent GRPO extensions [flowgrpo, dancegrpo] use GRPO to diffusion models by coupling the training loss with SDE samplers. Although significant progress has been made in RL for diffusion models, most of the work has focused on multi-step models. Here, we find that these algorithms are also adapted to the few-step model when carried out in conjunction with distribution matching distillation.
Figure 2: Overview of DMDR, which contains three key elements: (1) A DMD framework to optimize the generator by utilizing the gradient derived from an implicit distribution matching objective; (2) A RL branch to concurrently incorporate reward feedback from the reward model; (3) Two dynamic training strategies to facilitate more efficient and effective distillation during the initial phase.
3 Method
The pipeline of DMDR is shown in Figure 2, which contains three key elements: A DMD branch to optimize the generator using the gradient of an implicit distribution matching objective (Section 3.1); A RL branch to synchronously introduce the reward feedback from the reward model (Section 3.2); Two dynamic training strategies to achieve faster and better distillation in the initial phase (Section 3.3).
3.1 Preliminary
As our goal is to obtain a few-step model with performance beyond its multi-step teacher, and our method is built upon DMD. Here we provide a brief introduction to DMD:
Distribution Matching Distillation (DMD) [dmd] compresses a multi-step diffusion model (teacher) into a few-step generator (student) $G$ by minimizing the time-averaged approximate Kullback-Leibler (KL) divergence between the real distribution $p_{\text{real},t}$ and the synthetic distribution $p_{\text{fake},t}$. Note that $G$ is optimized via gradient descent, and this gradient admits a compact expression as the difference of two score functions:
$\nabla_{\!\theta}\,\mathcal{L}_{\text{dmd}}$ $=\mathbb{E}_{t}\!\left[\nabla_{\!\theta}\operatorname{KL}(p_{\text{fake},t}\|p_{\text{real},t})\right]$ (1)
$=-\mathbb{E}_{t}\!\left[\int\!\Bigl(s_{\text{real}}\bigl(F_{t})\bigr)-s_{\text{fake}}\bigl(F_{t}\bigr)\Bigr)\frac{\mathrm{d}G_{\theta}(z)}{\mathrm{d}\theta}\,\mathrm{d}z\right]\!,$
where $z\sim\mathcal{N}(0,\mathbf{I})$, $\theta$ denotes the parameters of $G$, and $F_{t}$ is the forward diffusion operator that injects noise at time $t$ to $G_{\theta}(z)$. The quantities $s_{\text{real}}$ and $s_{\text{fake}}$ are the score functions estimated by diffusion models $\mu_{\text{real}}$ and $\mu_{\text{fake}}$, respectively. During training, the $\mu_{\text{fake}}$ is updated via diffusion loss $\mathcal{L}_{\text{diff}}$ (for denoising-based models, the prediction target is noise, while for flow-based ones, it is velocity) on synthetic samples produced by the few-step generator. Moreover, as introduced in DMD2 [dmd2], $G_{\theta}(z)$ is a noisy synthetic image produced by the current $G$ running several steps, which is called backward simulation. The generator $G$ then denoises these simulated images, and the outputs are supervised with the above loss functions.
3.2 DMDR: Brings RL into DMD
It is worth noting that in DMD, the goal of the student generator is to align the output distribution as closely as possible with the distribution learned by the multi-step teacher. However, this also causes the ability of the student model to be capped by the teacher. Unlike the previous methods [dmd2, magic141] that introduce external real data to remedy this, we address the problem in an image-free manner by introducing the recently popular diffusion RL [refl, diffusiondpo, flowgrpo] in the DMD pipeline. Below, we explain how DMD and RL benefit from each other when training together.
RL Unlocks the Performance of DMD. One limitation of DMD is that the student model’s performance is inherently capped by the teacher model. By incorporating an RL objective, we allow the student model to receive supervision signals beyond merely mimicking the teacher, thus making it possible to surpass the teacher in an image-free training setup.
Figure 3: Illustration for mode seeking process.
Specifically, we integrate the reward models which provide reward signals indicating the specific characteristic (e.g, quality, text-alignment) of a generated image. During the joint training process, the RL component guides the student generator $G$ to produce samples that maximize these rewards. This has a direct impact on the DMD process: (1). Helping to jump out of the bad modes: The teacher’s distribution might contain modes that are less desirable or have lower perceived quality according to the reward model. RL actively discourages the student from generating samples within these low-reward regions, and “pulls” the student’s distribution towards regions in the image space that yield higher reward scores even if they are present in the teacher’s distribution. (2) Helping to mitigate zero forcing: In DMD, there’s a risk of “zero forcing,” where the student might fail to cover certain modes present in the teacher’s distribution, as we show the first row of Figure 3. By actively guiding the student towards high-reward modes, RL helps ensure that these important and preferred modes are reshaped, thus adequately covered during distillation, preventing their accidental omission as shown in the second row of Figure 3.
DMD Can Effectively Regularize RL. In addition to maximizing rewards, another important point in RL is the regularization, which helps to avoid rapid overfitting of the reward (reward hacking). Although this regularization takes on different forms in various methods (e.g, pretrain loss in ReFL [refl], KL divergence loss in Flow-GRPO [flowgrpo]), essentially, they all aim to let the training model be close to the reference model (always the one before RL). In the multi-step model, this is tractable because there is a pre-trained reference model. But in the few-step model, the situation is somewhat different. As shown in Figure 4, when RL is naively performed after distillation, the goal of regularization becomes to close the new model’s distribution $P_{\text{fake-new}}$ to the distilled few-step model’s $P_{\text{fake-ref}}$. However, as mentioned above, some modes are inevitably neglected after being distilled, thus $P_{\text{fake-ref}}$ cannot serve as an effective regular expression, as discovered by other works [hypersd] that these lead to quick hacking. However, if RL and DMD are carried out simultaneously, this problem will be easily solved. The DMD loss (Equation 1) continuously pulls the student generator’s distribution ($P_{\text{fake}}$) towards the multi-step teacher’s distribution ($P_{\text{real}}$). If viewed from the RL side, this also means that at every step of the joint training, the student is being regularized by the comprehensive and robust $P_{\text{real}}$ itself (the solid arrow “DMD + RL” in Figure 4), which reduces error accumulation in regularization and the risk of the reward hacking (see Figure 9).
Putting together, the final loss function can be written as:
$\mathcal{L}=\mathcal{L}_{\text{dmd}}+\mathcal{L}_{\text{rl}},$ (2)
where $\mathcal{L}_{\text{rl}}$ is the plug-and-play loss from the RL branch, the gradient can be differentiable type like ReFL, or policy type like GRPO, etc. In practice, we will multiply a coefficient if the two losses are not on the same scale.
3.3 A Dynamic Cold Start Stage for DMDR
While DMDR enables the student to surpass its teacher, the reward model often requires the generator to possess a basic generation capability to produce valid rewards [refl, srpo]. In practice, we observe that reward signals are noisy and unreliable in the early training phase. To this end, we conduct a cold start phase and design two dynamic training strategies based on vanilla DMD to accelerate and improve it.
Figure 4: Visualization of distribution matching directions.
Figure 5: Visual comparison between the teachers, selected competing methods [dmd2, hypersd, ladd], and ours. All images are generated using identical noise. Our model produces images with superior quality and prompt coherence. More comparisons are available in Appendix B.
Dynamic Distribution Guidance (DynaDG). In the initial stage of training, due to the significant disparity in image quality (the generator is unable to generate qualified images), the distribution of these generated samples ($P_{fake}$) is completely different from that of the training samples of the real score estimator ($P_{real}$), which finally leads to the inaccuracies in the KL divergence approximation as we show in Figure 6 (left). This substantial divergence comes from sparse or non-existent overlap between $P_{real}$ and $P_{fake}$, rendering the gradient $\nabla\log P_{real}-\nabla\log P_{fake}$ unreliable for effective optimization. Inspired by a recent study [magic141], we introduce a LoRA set [lora] and inject it not only into the fake score estimator to learn the rapidly evolving distribution of the few-step generator but also strategically into the real score estimator. The core idea behind our DynaDG is to dynamically adjust the distribution estimated in the real score branch. By injecting LoRA into the real score estimator with a smaller scale, we can effectively “pull” the perceived $P_{real}$ towards the nascent $P_{fake}$ distribution, creating a region of “more overlap” as depicted in Figure 6 (right). This increased overlap ensures that the estimated score difference, $\nabla\log P_{real}-\nabla\log P_{fake}$, consistently provides a “right direction estimation and optimization” signal, even when the distributions are initially far apart. Crucially, this LoRA scale for the real score estimator gradually weakens as training progresses. This mechanism allows the real score estimator to progressively revert to accurately representing the true $P_{real}$ as the generator improves and $P_{fake}$ naturally converges towards $P_{real}$.
Figure 6: Illustration for our Dynamic Distribution Guidance.
Dynamic Renoise Sampling (DynaRS). Another important operation in DMD is renoising, which renoises the samples generated by the generator for estimating the score (i.e., the distribution) associated with the specific noise level. The core idea of our DynaRS is to initially bias the sampling of renoise levels $t$ in $F_{t}$ towards higher values, and then gradually transition to a uniform sampling like DMD dose. This has two advantages for a faster and better cold start. First, when biasing to high noise, the forward operator makes noise largely occupy the image, so even bad samples produced by the generator in the initial stage lie in a similar regime for both real and fake estimators, thus improving score reliability and overlap. Second, the estimated score at a higher noise level is more about global structure [lsast, prospect], which can guide the generator to first capture global structure. As the capacity of the generator improves, the renoise bias is progressively reduced, allowing the model to learn finer details by sampling from a wider range of noise levels.
3.4 Distinction from Prior Few-step RL Works
Some prior works also try to introduce RL for the few-step diffusion model [hypersd, ladd, hyperbagel, hypernoise], but they mainly perform it after the distillation, and the problem mentioned by Hyper-SD [hypersd] of quickly deviating from the original domain still occurs. In this work, we illustrate the rationale for conducting the two simultaneously and how DMD loss can effectively regularize RL to mitigate the above problems. Meanwhile, although pioneer work like Diff-Instruct ++ [diffinstruct++] also introduces RL in the distillation process, our motivation and analysis are distinct, with more intuitive illustrations of the mutual benefits of the two, the verification of different diffusion models and RL algorithms at a large scale, as well as the dynamic cold start stage before incorporating RL.
4 Experiments
4.1 Experimental Setup
We provide a brief overview of our training and evaluation schemes here.
For training, we adopt denoising-based model (SDXL-Base [sdxl]) and flow-based models (SD3-Medium [sd3], SD3.5-Large [sd35]) for distillation, using prompts from t2i-2M [t2i-2M]. Additionally, we use ReFL [refl] with DFN-CLIP [dfnclip] and HPSv2.1 [hpsv2] as reward models by default. As for evaluation, we validate the effect of DMDR through extensive experiments. First, we use prompts sampled from a recent public dataset ShareGPT-4o-Image [sharegpt4o-image] to generate images and follow previous work [dmdx, tdm] that report CLIP Score [clipscore], Aesthetic Score [laion5b], Pick Score [pickscore], and Human Preference (HP) Score [hpsv2]. Next, in order to obtain a more comprehensive evaluation, we also compare our distilled models with their teachers on two popular benchmarks, DPG_Bench [DPG] and GenEval [geneval]. Finally, we conduct both qualitative and quantitative ablation experiments to further verify our method’s effectiveness.
Table 1: System-level comparison against state-of-the-art methods. denotes our reproduced results. Best performance is marked in Bold. Image-Free denotes whether the training requires external image data or data generated by the multi-step teacher.
[…]
SDXL-Base
Base-Model (CFG = 7.0) 25 50 34.7588 5.6480 22.1085 27.1477 -
LCM [lcm] (CFG = 8.0) 1 2 28.4664 5.1026 20.0603 17.6837 $\times$
Lightning [sdxl-lightning] 1 1 32.0283 5.6761 21.4868 26.3615 $\times$
DMD2 [dmd2] 1 1 34.3046 5.6238 21.7293 26.5986 $\times$
DMDR (ours) 1 1 35.4835 6.0483 22.5424 31.1442 $\checkmark$
DMD2 [dmd2] 4 4 34.5169 5.7043 22.1546 28.5655 $\times$
DMDR (ours) 4 4 35.2940 5.9857 22.6268 32.8678 $\checkmark$
SD3-Medium
Base-Model (CFG = 7.0) 25 50 34.9025 5.5942 22.1801 28.4021 -
Hyper-SD [hypersd] (CFG = 5.0) 4 8 32.0234 5.2489 20.2831 22.4544 $\times$
DMD2 [dmd2] 4 4 33.9421 5.6137 21.7688 27.3675 $\times$
Flash [flashdiff] 4 4 34.2634 5.6702 21.5921 26.6542 $\times$
TDM [tdm] 4 4 34.0301 5.6250 22.0010 27.7522 $\checkmark$
DMDR (ours) 4 4 34.9542 5.8462 22.3578 31.8979 $\checkmark$
SD3.5-Large
Base-Model (CFG = 3.5) 25 50 35.5509 5.7014 22.4856 28.8135 -
LADD [ladd] 4 4 35.0480 5.4514 22.2451 27.8470 $\times$
DMDR (ours) 4 4 35.8647 6.0284 22.8859 32.4724 $\checkmark$
Table 2: Quantitative comparison on DPG_Bench [DPG] of our four-step model against its multi-step teacher.
Model Overall Global Entity Attribute Relation Other
SDXL-Base
Teacher 74.65 83.27 82.43 80.91 86.76 80.41
Ours 76.44 83.70 82.54 83.66 84.75 83.49
SD3-Medium
Teacher 84.08 87.90 91.01 88.83 80.70 88.68
Ours 84.96 90.46 90.05 90.52 87.12 91.23
SD3.5-Large
Teacher 84.12 91.48 90.22 87.81 91.20 89.49
Ours 85.30 90.46 90.50 90.66 87.40 90.18
4.2 Main Results
We begin by assessing DMDR’s text-to-image generation performance using prompts sampled from ShareGPT-4o-Image [sharegpt4o-image], ensuring that there is no overlap with the prompts utilized during training. Table 4.1 presents a summary of the distillation results in comparison to other methods. Models distilled through our approach exhibit state-of-the-art (SOTA) capabilities in terms of prompt coherence and the generation of high-quality, aesthetically pleasing images. Notably, our method demonstrates effectiveness across various architectures (e.g., UNet, Transformer), model sizes (e.g., 2B, 8B), and paradigms (e.g., flow-based, denoising-based), highlighting the broad applicability and universality of the proposed approach. Additionally, we provide qualitative comparisons in Figure 5 and Appendix B, where our method produces images characterized by superior quality and enhanced prompt coherence.
To further verify our method’s effectiveness and provide a more comprehensive comparison between the few-step distillation model and the multi-step teacher model, we conduct the comparison on two commonly used benchmarks [DPG, geneval], Table 2 shows our four-step model consistently outperforms its multi-step teacher on DPG_Bench across all base models, lifting the overall score from 74.65 to 76.44 for SDXL-Base, 84.08 to 84.96 for SD3-Medium, and 84.12 to 85.30 for SD3.5-Large. Meanwhile, on GenEval (Table 3) the same trend holds: overall scores improve for SDXL (0.55 → 0.56), SD3 (0.62 → 0.64), and SD3.5 (0.71 → 0.72). These results indicate that DMDR has successfully freed the student model from the constraints of the multi-step teacher model and stimulated its capabilities during the distillation process.
Table 3: Quantitative comparison on GenEval [geneval] of our four-step model against its multi-step teacher.
Model Overall Single Two Count. Colors Pos. Attr.
SDXL-Base
Teacher 0.55 0.98 0.74 0.39 0.85 0.15 0.23
Ours 0.56 0.99 0.76 0.42 0.84 0.11 0.24
SD3-Medium
Teacher 0.62 0.98 0.74 0.63 0.67 0.34 0.36
Ours 0.64 0.99 0.84 0.54 0.81 0.25 0.44
SD3.5-Large
Teacher 0.71 0.98 0.89 0.73 0.83 0.34 0.47
Ours 0.72 0.99 0.93 0.68 0.80 0.30 0.59
Figure 7: Ablation examples of two dynamic training strategies. Better to zoom in to check the effect.
4.3 Ablation Study
We investigate our method’s properties below. We use SD3-Medium with $512\times 512$ resolution by default for a fast ablation.
Effect of Dynamic Cold Start Stage. As shown in Table 4.3, our dynamic cold start strategies significantly enhance the initial training phase. Firstly, introducing DynaDG leads to a notable improvement. By dynamically adjusting the $P_{real}$ to create more overlap with the nascent $P_{fake}$ distribution, DynaDG provides more reliable gradient signals. Quantitatively, this translates to a CLIP Score increase from 31.56 to 32.79 (+1.23) and an HP Score increase from 27.01 to 27.57 (+0.56) at 1000 iterations. Similar gains are observed at 2000 iterations. Further incorporating DynaRS, which initially biases renoise sampling towards higher noise levels, guiding the generator to learn global structures before progressively shifting to finer details, further boosts performance. To highlight the importance of the dynamic aspect of our strategies, we evaluate a variant denoted as -Dynamic. This variant removes the dynamic scaling of LoRA in DynaDG (e.g., using a fixed LoRA scale) and employs biased renoise sampling in the whole training process. While it still outperforms vanilla_DMD, it consistently underperforms the full dynamic approaches. This clearly indicates that the adaptive, progressive nature of our DynaDG and DynaRS mechanisms is crucial for maximizing performance during the cold start phase. Qualitative results presented in Figure 7 further corroborate these findings, showing a faster global structure building when our dynamic cold start strategies are employed.
Table 4: Ablation on our dynamic cold start stage.
Method 1000 iter 2000 iter
CLIP Score HP Score CLIP Score HP Score
vanilla_DMD 31.5608 27.0051 32.5492 28.4541
   +DynaDG 32.7905 27.5723 33.5045 29.0363
   +DynaRS 33.1657 27.8276 33.6432 29.1157
   -Dynamic 32.6562 27.2641 33.0245 28.8459
Effect of DMDR with Different RL Algorithms. We choose three representative RL methods [refl, diffusiondpo, flowgrpo] to illustrate the efficacy of combining distillation with various RL techniques in DMDR, the results are provided in Table 4.3. Firstly, we observe that the initial model state after our dynamic cold start stage already provides a good baseline performance. Training with “w/ only Distill.”, as we illustrated above, this would be inherently capped by the teacher’s distribution, resulting in a very minor improvement over the initial state. Next, when considering introducing different RL algorithms (the regularization originally used in these methods has also been added), although these show improvements in some metrics, they do not consistently surpass the “only Distill” baseline across all scores. Moreover, the reward hacking occurs (we will analyze it specifically in the next part). Crucially, the rows corresponding to “w/ Distill. + RL” consistently demonstrates superior performance across all evaluated metrics against the counterparts. Figure 8 also shows superior visualization results when DMD and RL are conducted together. These validate our core insight that RL unlocks the performance potential of DMD, allowing the student model to surpass the teacher’s inherent limitations by optimizing for external reward signals. Simultaneously, DMD effectively regularizes the RL process, preventing reward hacking and ensuring that the student’s distribution remains robust and comprehensive, as evidenced by the consistent improvements across diverse metrics with different RL algorithms.
Table 5: Ablation on RL and DMD. We init the model after cold start for 2K iteration then continue training in different ways.
Method CLIP Aesthetic Pick HP
Score Score Score Score
init 33.6432 5.6124 21.0489 29.1157
w/ only Distill. 33.6738 5.6248 21.6376 29.1389
w/ only RL (ReFL) 33.1897 5.8841 22.3008 31.2714
w/ Distill. + RL (ReFL) 34.6249 6.1813 22.7578 32.8979
w/ only RL (DPO) 33.9554 5.7893 21.7643 29.5819
w/ Distill. + RL (DPO) 33.9632 5.9710 21.9865 30.5994
w/ only RL (GRPO) 33.9095 5.7876 21.7956 29.4429
w/ Distill. + RL (GRPO) 34.0055 5.8256 22.0120 30.6248
Figure 8: Ablation examples of different RL algorithms with DMD. Better to zoom in to check the effect.
Furthermore, an interesting observation is that using the ReFL style RL method achieves overall better performance, and our analysis is that: traditional ReFL for multi-step models ignores earlier sampling steps and only trains the last few steps before the output image [refl] because the large computational graph generated by the gradient of the reward model and the multi-step denoising process would result in a huge memory consumption. Subsequent works [srpo, deepreward] mainly focus on saving forward noise trajectory and sampling a subset of steps to optimize to remedy this dilemma, although this can allow the gradient to return to the early denoising step, it is also prone to error accumulation. On the contrary, in our few-step model, the model is trained to predict a clean image at each step’s state [dmd2], which enables the gradient of ReFL to directly return to the initial state, thus obtaining an effective optimization.
Figure 9: Reward curves comparison during the RL stage. Better to zoom in to check the difference.
Effect of DMDR to Mitigate Reward Hacking. We follow RewardDance [rewarddance] to present the reward curves during training illustrated in Figure 9. It is evident that employing DMD loss as a form of regularization (right) demonstrates significantly greater variance throughout the training process compared to the traditional approach (left). This increased diversity in reward signals suggests that models trained using our method exhibit enhanced robustness against reward hacking. This assertion is further supported by Figure 8, which shows that in the absence of DMD loss, utilizing ReFL results in a rasterized appearance of images, indicating the occurrence of reward hacking.
5 Limitation and Future Work
We discuss some limitations we realized during the development of DMDR. We hope these will serve as a catalyst for innovative ideas and advancements in future research.
Quality vs. Diversity. One notable thing is the trade-off between quality and diversity. Although models trained from our method show superior quality and prompt coherence, the diversity is reduced like shown in Table 6. We believe that this is a common feature of score distillation like DMD and RL, that is, sacrificing uncertainty in sampling in exchange for stable and desired results, not a limitation unique to our approach. The reduced diversity caused by DMD, we believe, will also have some impact on some online RL methods that require a group of samples for ranking then optimization. This is also reflected in the slightly poorer improvement of GRPO and DPO compared to ReFL in the Table 4.3. Further investigation into finding the optimal balance between quality and diversity is left for future work.
Table 6: Diversity evaluation on ShareGPT-4o-Image [sharegpt4o-image].
DMD DMD2 Ours Teacher
[…]
Scaling of Reward Model. Another notable thing is the scaling of the reward model in DMDR. Our experimental results demonstrate that our method enables the few-step model to be able to surpass its teacher model. However, as indicated in the Table 2 and 3, certain metrics exhibit a decline after the distillation. We attribute this degradation primarily to inherent limitations and potential biases present in the open-source reward model we employed. Notably, RewardDance [rewarddance] has shown that scaling up the reward model yields consistent improvements for RL in multi-step diffusion models. We posit that similar benefits would extend to few-step models with our method, but require more computation resources, which we leave for future work.
6 Conclusion
This work introduces DMDR, which combines Distribution Matching Distillation (DMD) with reinforcement learning (RL) in one framework. We show that DMD and RL can be trained simultaneously, with RL enabling the student model to surpass the teacher and DMD loss regularizing RL to prevent reward hacking. We also propose dynamic training strategies (DynaDG and DynaRS) to enhance the initial distillation process. Experiments demonstrate state-of-the-art performance among few-step methods and overall superior performance against multi-step teachers.
Appendix
Appendix A More Related Works
Trajectory-Based Distillation. Trajectory-based distillation [lcm, hypersd, rcm, piflow], which typically aims to simulate teacher ODE trajectories on the instance level. Early work [kdgen] regresses the teacher’s ODE integral in one step, producing blurry $\ell_{2}$$x_{0}$ estimates, and thus suffers from degraded quality. Progressive distillation [hypersd, shortcut-model] mitigates this by a multi-stage pipeline that enlarges the student’s step size and halves its NFE each stage by distilling the prior stage’s trajectory into fewer steps, However, this not only makes the training a multi-stage process (e.g., NFE from 50 to 25 then to 10, etc.), but also leads to cumulative errors. Meanwhile, consistency models [lcm, scm] aim to learn a consistency function that maps the point at an arbitrary time t on the teacher’s PF-ODE trajectory to the initial point. However, the student model must be constructed implicitly using either inaccurate finite differences or expensive Jacobian–vector products (JVPs) and the quality is still limited due to the accumulation of errors into the integrated state, which limits the application in large-scale scenarios [rcm, piflow].
Adversarial Distillation. Adversarial distillation [add, apt, ladd, apt2] can be regarded as another form of distribution matching distillation. The difference is that adversarial distillation aim to estimate the distribution of both the student model and the real data or the teacher model through a discriminator model [gan] to match the distribution. Representative works like Adversarial Diffusion Distillation (ADD) [add] force the few-step student diffusion model to fool a discriminator which is trained to distinguish the generated samples from real images. Follow-up works [apt, apt2] applied this idea to other fields. However, such a training method can cause many instabilities, similar to traditional GANs [gan, improvedgan], hence requires a lot of tricks to stabilize training [animatediff-l, r1reg, train-gan].
Other Distillation Works. Other distillation works mainly focus on combining the distillation at the distribution level and the trajectory level [sana-sprint, rcm, tdm, pose]. For example, SANA-Sprint [sana-sprint] combines the ideas of LADD [ladd] and sCM [scm] and achieves fast convergence and high fidelity generation while retaining the alignment advantages of sCMs. TDM [tdm] and rCM [rcm] combine DMD [dmd] and LCM [lcm] together to remedy the quality issues of LCM.
Appendix B Additional Visual Comparison Results
In this section, we provide additional visual comparisons for our 4-step distilled models against teacher models and other methods. The results are shown in Figure 10, Figure 11:
Figure 10: Additional visual comparison in SDXL-Base. All images are generated using identical noise and text prompts. Our model produces images with superior quality and prompt coherence.
Figure 11: Additional visual comparison in SD3-Medium. All images are generated using identical noise and text prompts. Our model produces images with superior quality and prompt coherence.
Figure 12: Additional visual comparison in SD3.5-Large. All images are generated using identical noise and text prompts. Our model produces images with superior quality and prompt coherence.

### v4

@@
 Abstract
-Distribution Matching Distillation (DMD) distills a pre-trained multi-step diffusion model to a few-step one to improve inference efficiency.
-However, the performance of the latter is often capped by the former.
-To circumvent this dilemma, we propose DMDR, a novel framework that combines Reinforcement Learning (RL) techniques into the distillation process.
-We show that for the RL of the few-step generator, the DMD loss itself is a more effective regularization compared to the traditional ones.
-In turn, RL can help to guide the mode coverage process in DMD more effectively.
-These allow us to unlock the capacity of the few-step generator by conducting distillation and RL simultaneously.
-Meanwhile, we design the dynamic distribution guidance and dynamic renoise sampling training strategies to improve the initial distillation process.
-The experiments demonstrate that DMDR can achieve leading visual quality, prompt coherence among few-step methods, and even exhibit performance that exceeds the multi-step teacher.
+Distribution Matching Distillation (DMD) facilitates efficient inference by distilling multi-step diffusion models into few-step variants.
+Concurrently, Reinforcement Learning (RL) has emerged as a vital tool for aligning generative models with human preferences.
+While both represent critical post-training stages for large-scale diffusion models, existing studies typically treat them as independent, sequential processes, leaving a systematic framework for their unification largely unexplored.
+In this work, we demonstrate that jointly optimizing these two objectives yields mutual benefits: RL enables more preference-aware and controllable distillation rather than uniformly compressing the full data distribution, while DMD serves as an effective regularizer to mitigate reward hacking during RL training.
+Building on these insights, we propose DMDR, a unified framework that incorporates Reward-Tilted Distribution Matching optimization alongside two dynamic distillation training strategies in the initial stage, followed by the joint DMD and RL optimization in the second stage.
+Extensive experiments demonstrate that DMDR achieves state-of-the-art visual quality and prompt adherence among few-step generation methods, even surpassing the performance of its multi-step teacher model.
 1 Introduction
-Diffusion models have achieved unparalleled quality in visual generation tasks [dit, sd3, flux, lumina2, sdxl].
-Yet their sampling process typically demands tens of iterative denoising steps, each involving a full forward pass through a large neural network, which makes high-resolution text-to-image synthesis costly both in time and computation.
-To accelerate sampling speed, a surge of work [diffinstruct, dmd, lcm, hypersd, dmd2, flashdiff] has sought to distill the original diffusion model into a generator that needs only a handful of steps (e.g, 4).
-These approaches can be roughly divided into two types: distillation from the trajectory and distillation from the distribution.
-The former method operates on the instance level, aiming to simulate the original teacher diffusion generation trajectories in as few steps as possible [lcm, piflow].
-While the latter works at the distribution level, striving to match the output distribution of the few-step student generator to the original multi-step one’s [diffinstruct, dmd].
-In this study, we focus on distribution matching distillation (DMD) as it has been verified in large-scale settings [dmd2, dmdx] and has already been used in some leading products (e.g, Seedream 4.0 [seedream4]).
-However, one notable thing is that in DMD, the student model aims to match the distribution of the teacher model, which also means that the ability of the student model is often limited by the teacher.
-To remedy this, most works (e.g., DMD2 [dmd2]) introduce an additional GAN branch for guidance [gan].
-However, training with GAN not only renders the training process dependent on external high-quality visual data, but also introduces training instability inherent in GAN [wgan, improvedgan].
-To this end, we propose DMDR, which combines distribution matching distillation (DMD) with recent powerful diffusion reinforcement learning (RL) [refl, flowgrpo, diffusiondpo, ddpo].
-In a nutshell, DMDR enables the student model to obtain supervision signals beyond the teacher model without relying on external image data, and makes it possible for the student to surpass the teacher.
-Contrary to the prevailing belief that RL should be conducted after obtaining the few-step generator, we perform RL concurrently with the distillation process.
-We show that such a combination is mutually beneficial: (1) As for DMD, RL can help to reshape some existing low-probability modes by guiding them toward higher reward regions, ensuring these modes are covered during the distillation process, thereby mitigating the risk of zero forcing.
-(2) As for RL, DMD loss itself can serve as a more effective regularization than its traditional counterparts, thereby mitigating the risk of reward hacking.
-Meanwhile, since obtaining valid reward scores requires the few-step model to possess a basic generation capability already, we design two dynamic training strategies to facilitate a faster and better “cold start” for DMDR: (1) DynaDG, which strengthens the initial distributional overlap between the teacher and student model; (2) DynaRS, which enhances the student’s global modeling of the teacher’s distribution during the early training phase.
-Taken together, our experimental results show that our method leads to state-of-the-art few-step generative models that can outperform their teacher, using as few as 4 sampling steps.
+Diffusion models pre-trained on large-scale datasets have achieved remarkable performance in visual generation tasks [dit, sd3, flux, zimage, sdxl].
+However, their sampling process typically requires numerous iterative denoising steps [ddim, ddpm, flow-matching], each involving a full forward pass through a large neural network.
+This requirement makes high-resolution text-to-image synthesis computationally demanding.
+Furthermore, base models often fail to produce images that align closely with human preferences and aesthetic criteria [pickscore, hpsv2, hpsv3].
+To enhance practical utility, researchers typically employ two post-training stages.
+First, step-distillation [diffinstruct, dmd, lcm, hypersd, dmd2, flashdiff], which compresses the original model into a generator capable of few-step sampling (e.g., 4 steps) for efficient sampling.
+Among such methods, Distribution Matching Distillation (DMD) [dmd] is widely recognized for its effectiveness in large-scale scenarios [dmd2, dmdx, twinflow, senseflow] and has been adopted in prominent industrial applications [seedream4, zimage, flux-2].
+Second, reinforcement learning (RL) [refl, flow-dpo, flowgrpo, ddpo, dancegrpo, srpo, diffusionnft, diffusiondpo] aligns the model with human preferences to improve aesthetic quality.
+Despite progress in both domains, most existing studies address these objectives in isolation.
+Early efforts to incorporate RL into the distillation process were largely confined to a two-stage paradigm [diffinstruct++, hypersd, pso, hypernoise] (exploring how to conduct RL on an already distilled model).
+Consequently, there remains a critical gap in the field: a systematic framework to explore the unification and potential synergy between these two “temporarily separated” stages.
+To bridge this gap, we propose DMDR, a unified framework that seamlessly integrates DMD and RL into a synergistic training pipeline.
+Our core insight is that concurrent optimization yields mutual benefits: (1) Preference-Aware Distillation: RL steers the distillation process toward high-reward regions, breaking the ”performance ceiling” inherent in purely mimicking a teacher; (2) Distributional Regularization: The DMD objective serves as a robust regularizer for RL, mitigating reward hacking by maintaining proximity to the teacher’s manifold.
+Specifically, DMDR operates in two stages.
+In the initial stage, we implement Reward-Tilted Distribution Matching (RT-DM), which distills toward a reward-tilted teacher distribution to prioritize human-preferred regions.
+Concurrently, we introduce Dynamic Distribution Guidance (DynaDG) and Dynamic Renoise Sampling (DynaRS) to stabilize the ”cold-start” phase through annealed distributional overlap maximization.
+In the second stage, we transition to a joint optimization paradigm.
+In this phase, the RL objective drives aggressive preference alignment, while the DMD mechanism serves as a robust regularizer to prevent reward hacking.
+Our experimental results show that our method leads to state-of-the-art few-step generative models.
 Moreover, our method is not only compatible with different models (e.g., flow-based, denoising-based) but also with various RL algorithms (e.g., ReFL, DPO, GRPO).
@@
   • We propose DMDR, which shows that DMD and RL can be trained simultaneously with mutual benefits.
-  • We design a dynamic cold start stage for faster and better distillation in the initial phase.
-  • We validate our method on various models and RL algorithms with all of which achieve excellent performance.
-Figure 1: More Images generated by Z-Image-Turbo [zimage] distilled through our DMDR.
-2 Related Work
-We highlight key related studies here, and leave the discussions of others in Appendix A.
+  • We design a Reward-Tilted Distribution Matching and two dynamic distillation training strategies for better integrating RL and distillation in the initial phase.
+  • We validate our method on various models and RL algorithms, all of which achieve excellent performance.
+2 Method
+2.1 Preliminary
 Distribution matching distillation.
-Distribution Matching Distillation (DMD) [dmd] is the first work to successfully apply the principle score-based distillation [score-base-diff, prolificdreamer, dreamfusion] to large-scale text-to-image models.
-Intuitively, it strives to ensure that any sample realized by the student at a given noise level occurs with exactly the same probability as it would under the teacher’s distribution, thereby preserving the multi-step generative priors in the few-step model.
-From then on, a lot of follow-up work emerged [dmd2, dmdx, tdm, sd3.5flash, f-distill], for example, DMD2 [dmd2] employs a discriminator to align the student model distribution with a specific target distribution like GAN [gan] does; TDM [tdm] incorporates DMD loss in the sampling process of the student model for better alignment; ff–distill [f-distill] replaces the original reverse Kullback–Leibler (KL) divergence to the proposed $f$-divergence for covering different divergences with different properties.
-Our work also starts with DMD, but we don’t focus on how to better “imitate” the teacher.
-Instead, we aim to incorporate reinforcement learning in the distillation process to enable the student to outperform the teacher.
-Reinforce Learning for Diffusion Models.
-Reinforce learning helps to align diffusion models to human preferences by training a reward model and using it to guide generation [ddpo, flowgrpo, refl, srpo, diffusiondpo, dancegrpo].
-There are many algorithms to conduct RL, for example, DDPO [ddpo] adapts PPO via image log-likelihoods; ReFL [refl] bypasses likelihoods by optimizing outputs with frozen-reward gradients; Diffusion-DPO [diffusiondpo] adapts DPO to diffusion for paired human preference data; and recent GRPO extensions [flowgrpo, dancegrpo] use GRPO to diffusion models by coupling the training loss with SDE samplers.
-Although significant progress has been made in RL for diffusion models, most of the work has focused on multi-step models.
-Here, we find that these algorithms are also adapted to the few-step model when carried out in conjunction with distribution matching distillation.
-Figure 2: Overview of DMDR, which contains three key elements: (1) A DMD framework to optimize the generator by utilizing the gradient derived from an implicit distribution matching objective; (2) A RL branch to concurrently incorporate reward feedback from the reward model; (3) Two dynamic training strategies to facilitate more efficient and effective distillation during the initial phase.
-3 Method
-The pipeline of DMDR is shown in Figure 2, which contains three key elements: A DMD branch to optimize the generator using the gradient of an implicit distribution matching objective (Section 3.1); A RL branch to synchronously introduce the reward feedback from the reward model (Section 3.2); Two dynamic training strategies to achieve faster and better distillation in the initial phase (Section 3.3).
-3.1 Preliminary
-As our goal is to obtain a few-step model with performance beyond its multi-step teacher, and our method is built upon DMD. Here we provide a brief introduction to DMD:
-Distribution Matching Distillation (DMD) [dmd] compresses a multi-step diffusion model (teacher) into a few-step generator (student) $G$ by minimizing the time-averaged approximate Kullback-Leibler (KL) divergence between the real distribution $p_{\text{real},t}$ and the synthetic distribution $p_{\text{fake},t}$.
+DMD [dmd] compresses a multi-step diffusion model (teacher) into a few-step generator (student) $G$ by minimizing the time-averaged approximate Kullback-Leibler (KL) divergence between the real distribution $p_{\text{real},t}$ and the synthetic distribution $p_{\text{fake},t}$.
 Note that $G$ is optimized via gradient descent, and this gradient admits a compact expression as the difference of two score functions:
@@
 The generator $G$ then denoises these simulated images, and the outputs are supervised with the above loss functions.
-3.2 DMDR: Brings RL into DMD
-It is worth noting that in DMD, the goal of the student generator is to align the output distribution as closely as possible with the distribution learned by the multi-step teacher.
-However, this also causes the ability of the student model to be capped by the teacher.
-Unlike the previous methods [dmd2, magic141] that introduce external real data to remedy this, we address the problem in an image-free manner by introducing the recently popular diffusion RL [refl, diffusiondpo, flowgrpo] in the DMD pipeline.
-Below, we explain how DMD and RL benefit from each other when training together.
-RL Unlocks the Performance of DMD.
-One limitation of DMD is that the student model’s performance is inherently capped by the teacher model.
-By incorporating an RL objective, we allow the student model to receive supervision signals beyond merely mimicking the teacher, thus making it possible to surpass the teacher in an image-free training setup.
-Figure 3: Illustration for mode seeking process.
-Specifically, we integrate the reward models which provide reward signals indicating the specific characteristic (e.g, quality, text-alignment) of a generated image.
-During the joint training process, the RL component guides the student generator $G$ to produce samples that maximize these rewards.
-This has a direct impact on the DMD process: (1).
-Helping to jump out of the bad modes: The teacher’s distribution might contain modes that are less desirable or have lower perceived quality according to the reward model.
-RL actively discourages the student from generating samples within these low-reward regions, and “pulls” the student’s distribution towards regions in the image space that yield higher reward scores even if they are present in the teacher’s distribution.
-(2) Helping to mitigate zero forcing: In DMD, there’s a risk of “zero forcing,” where the student might fail to cover certain modes present in the teacher’s distribution, as we show the first row of Figure 3.
-By actively guiding the student towards high-reward modes, RL helps ensure that these important and preferred modes are reshaped, thus adequately covered during distillation, preventing their accidental omission as shown in the second row of Figure 3.
-DMD Can Effectively Regularize RL.
-In addition to maximizing rewards, another important point in RL is the regularization, which helps to avoid rapid overfitting of the reward (reward hacking).
-Although this regularization takes on different forms in various methods (e.g, pretrain loss in ReFL [refl], KL divergence loss in Flow-GRPO [flowgrpo]), essentially, they all aim to let the training model be close to the reference model (always the one before RL).
-In the multi-step model, this is tractable because there is a pre-trained reference model.
-But in the few-step model, the situation is somewhat different.
-As shown in Figure 4, when RL is naively performed after distillation, the goal of regularization becomes to close the new model’s distribution $P_{\text{fake-new}}$ to the distilled few-step model’s $P_{\text{fake-ref}}$.
-However, as mentioned above, some modes are inevitably neglected after being distilled, thus $P_{\text{fake-ref}}$ cannot serve as an effective regular expression, as discovered by other works [hypersd] that these lead to quick hacking.
-However, if RL and DMD are carried out simultaneously, this problem will be easily solved.
-The DMD loss (Equation 1) continuously pulls the student generator’s distribution ($P_{\text{fake}}$) towards the multi-step teacher’s distribution ($P_{\text{real}}$).
-If viewed from the RL side, this also means that at every step of the joint training, the student is being regularized by the comprehensive and robust $P_{\text{real}}$ itself (the solid arrow “DMD + RL” in Figure 4), which reduces error accumulation in regularization and the risk of the reward hacking (see Figure 9).
-Putting together, the final loss function can be written as:
-$\mathcal{L}=\mathcal{L}_{\text{dmd}}+\mathcal{L}_{\text{rl}},$ (2)
-where $\mathcal{L}_{\text{rl}}$ is the plug-and-play loss from the RL branch, the gradient can be differentiable type like ReFL, or policy type like GRPO, etc.
-In practice, we will multiply a coefficient if the two losses are not on the same scale.
-3.3 A Dynamic Cold Start Stage for DMDR
-While DMDR enables the student to surpass its teacher, the reward model often requires the generator to possess a basic generation capability to produce valid rewards [refl, srpo].
-In practice, we observe that reward signals are noisy and unreliable in the early training phase.
-To this end, we conduct a cold start phase and design two dynamic training strategies based on vanilla DMD to accelerate and improve it.
-Figure 4: Visualization of distribution matching directions.
-Figure 5: Visual comparison between the teachers, selected competing methods [dmd2, hypersd, ladd], and ours.
-All images are generated using identical noise.
-Our model produces images with superior quality and prompt coherence.
-More comparisons are available in Appendix B.
+Reinforce learning for diffusion models.
+Although there are various RL algorithms [refl, flowgrpo, diffusiondpo], All their goal is to optimize the diffusion model to generate the samples that can maximize a score given by reward models [hpsv2, pickscore, clipscore].
+This score can be expressed as aesthetic, texture, prompt following, and so on.
+At the same time, a regularization term is often added to balance the trade-off between reward maximization and the deviation from the original model [rlhf].
+Thus, the loss can be formulated as follows:
+$\mathcal{L}_{\text{rl}}=\mathbb{E}\left[r(\bm{x}_{0})-\operatorname{KL}(p_{new}\|p_{\text{ref}})\right],$ (2)
+where $r$ is the reward function, $p_{new}$ and $p_{ref}$ is the distribution of the trainable model and the reference (this can be a pre-train data distribution [refl] or the original model’s [flowgrpo, diffusiondpo, diffusionnft]), $\bm{x}_{0}$ is the clean image that is generated by the training generation model.
+2.2 Problem Formulation and Pipeline Overview.
+It is noted that DMD effectively reduces the sampling steps of the diffusion model, while RL enables the diffusion model to generate instruction-aligned and human-preferred images.
+Both of these are crucial for the practical application of the diffusion model.
+However, at present, the majority of the work [dmdx, dmd2, magicdistillation, flowgrpo, diffusionnft, dancegrpo] only focuses on studying and improving one aspect, without exploring the synergy between the two.
+And only a small portion of the works [pso, flash-dmd, diffinstruct++, hypersd, hypernoise] attempt to introduce RL into the distillation process, but they still focus on how to conduct RL on a distilled model.
+This raises a question: Can these two be combined into one stage and benefit each other?
+We address this challenge question by proposing DMDR, as shown in Figure 1, the training process of DMDR is divided into two operational stages: Stage 1: RT-DM with Dynamic Distillation.
+(Section 2.4 and Section 2.5) Stage 2: Joint RL + DMD Optimization(Section 2.3).
+The following sections will provide the detailed analysis and illustration.
+Figure 1: Overview of DMDR.
+It follows a two-stage training paradigm: Stage 1 (Reward-Tilted Distribution Matching) employs reward-weighted distillation to integrate preference signals into the early distillation phase, stabilized by Dynamic Distribution Guidance (DynaDG) and Dynamic Renoise Sampling (DynaRS).
+Stage 2 (Joint RL + DMD) performs direct reward maximization regularized by the DMD loss.
+2.3 Bring DMD and RL Together with Benefit
+We now show formally and empirically that DMD operates as a superior regularization mechanism that subsumes the traditional KL penalty in RL for few-step model, yielding robust joint optimization.
+Figure 2: Visualization of the difference of optimization directions.
+From the perspective of RL, standard preference optimization requires a penalty (e.g., $\operatorname{KL}(p_{new}\|p_{\text{ref}})$) to prevent reward hacking [refl, flowgrpo].
+In few-step models, sequentially applying RL after distillation uses the already-distilled distribution $P_{\text{fake-ref}}$ as the regularizer (Fig.
+2).
+Because distillation inherently loses some modes, $P_{\text{fake-ref}}$ acts as an impoverished anchor, leading to rapid overfitting and mode collapse [hypersd].
+More fundamentally, instead of viewing DMD and RL as two disparate losses added heuristically, we reframe the joint objective as maximizing reward under a distribution-level structural constraint.
+As the joint gradient can be decomposed into two orthogonal objectives [rlhf, score-base-diff, ppo]: a ”reward ascent” direction that increases human preference, and a ”manifold projection” term that corrects off-manifold displacement by pulling the student back toward the teacher’s regions.The DMD gradient $\nabla_{\!\theta}\mathcal{L}_{\text{dmd}}$ corresponds exactly to the maximum likelihood update projecting the student output distribution toward the learned data distribution of the multi-step teacher $P_{\text{real}}$.
+When combined linearly with the reward gradient $\nabla_{\!\theta}\mathcal{L}_{\text{rl}}$, any displacement induced by $r$ that exits the teacher’s high-density manifold generates a large discrepancy $s_{\text{real}}-s_{\text{fake}}$.
+The $\mathcal{L}_{\text{dmd}}$ gradient thus projects the reward ascent strictly along the teacher’s support.
+Simultaneously, from the perspective of distillation, pure DMD forces the student to uniformly mimic the teacher, implicitly establishing a ”performance ceiling” capped by distilling $P_{\text{real}}$ equally.
+Incorporating the RL objective addresses this fundamental limitation: RL steers the distillation toward high-reward regions, allowing the student to surpass the teacher while the DMD regularizer ensures semantic integrity is preserved.
+Figure 3: Verification of the joint optimization synergy.
+Directly conduct RL (red) on the distilled model exhibits high training rewards but suffers from reward hacking and semantic collapse (low DPG score).
+In contrast, the joint DMD+RL paradigm (green) effectively boosts the performance (better than the DMD baseline (blue) in all metric) while preventing reward hacking.
+Better to zoom in to check the effect.
+To validate the analysis, we conduct a comparative study using SD3-M [sd3] as the base model, with rewards provided by HPS v2.1 [hpsv2] and ReFL-based RL [refl] for training.
+We compare three paradigms: (i) DMD Only (Blue line); (ii) DMD then RL (Sequential: distillation followed by RL in red line); and (iii) DMD then DMD+RL (Joint approach: distillation followed by joint optimization in green line).
+We evaluate these models across training/test reward curves and the DPG_Bench score [dpg].
+As shown in Fig.
+3(a), when RL is introduced (after 1000 iterations), the sequential approach achieves the highest training reward.
+However, this gain is obtained through hacking score.
+Fig.
+3(b) and (c) reveal that this it suffers from severe reward hacking and catastrophic forgetting.
+While its training reward climbs, its test reward plateaus, and its DPG score collapses.
+This confirms that without the strong distributional anchor of the teacher, the few-step model quickly drifts into narrow, high-reward modes that sacrifice semantic integrity [hypersd].
+In contrast, the joint optimization demonstrates a clear synergistic effect.
+First, it effectively breaks the ”performance ceiling” of the teacher; the test reward significantly surpasses the DMD-only baseline.
+Second, by treating the DMD loss as a persistent structural regularizer, the model keeps improving DPG score while increasing aesthetic quality.
+This indicates that the DMD objective prevents the RL process from hacking, while the RL objective guides the distillation toward more human-preferred regions.
+Table 1: Comparison of different training strategies.
+The joint approach avoids the ”distillation gap” of ”RL then distillation” pipeline, achieving the highest preference score with $\sim 10\times$ rollout speedup.
+Method HPS DPG Time
+SD3-M, resolution $\times1024$, local batchsize 16 on H100
+Multi-Step RL 31.06 86.43 9.24s $\sim$ 10.57s
+Multi-Step RL then Distillation 30.42 85.92 -
+Few-Step Joint RL and DMD 31.42 86.08 0.43s $\sim$ 1.21s
+Based on the empirical synergy observed above, we formally define the optimization objective. The total loss is formulated as:
+$\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{dmd}}+\lambda_{\text{rl}}\mathcal{L}_{\text{rl}},$ (3)
+where $\mathcal{L}_{\text{dmd}}$ is the original distribution matching loss with the gradient formulated in Eq.
+1, and $\mathcal{L}_{\text{rl}}$ is the plug-and-play reward maximization loss from the RL branch (Eq.
+1), the gradient can be differentiable type like ReFL, or policy type like GRPO, etc.
+The hyperparameter $\lambda_{\text{rl}}$ is a balancing coefficient used to trade off the two losses.
+2.4 Reward-Tilted Distribution Matching for Integrating RL into the Early Phase of DMD
+While we have demonstrated that DMD and RL can be trained jointly with mutual benefit, we still rely on a two-stage paradigm that incorporates RL only after an initial ”cold start” distillation phase [hypersd, diffinstruct++].
+This requirement stems from the observation that reward models typically necessitate a baseline generation capability to produce reliable feedback [refl, srpo].
+Specifically, in DMD frameworks with flow matching model, the image used for gradient computation is defined as $\bm{x}_{t}-(0-t)G_{\theta}(\bm{x}_{t})$, where $\bm{x}_{t}$ is iteratively sampled.
+Notably, t can be any training time-step; thus, when training a 4-step model, the clean image can be generated in only one to four steps.
+In the early training phase, however, such few-step sampling often fails to produce coherent images, rendering reward signals noisy and unreliable for direct optimization.
+Figure 4: Effectiveness of RT-DM and two-stage training strategy.
+In Stage 1, the RT-DM paradigm (red/green) significantly accelerates preference alignment compared to the DMD baseline (blue).
+In Stage 2, while ”RT-DM only” (red) eventually saturates due to the constraints of the teacher’s manifold, the transition to joint DMD+RL (green) successfully breaks this performance plateau by directly optimizing the reward score.
+This complete two-stage approach achieves the highest test rewards and DPG scores.
+Better to zoom in to check the effect.
+To address this, we propose a Reward-Tilted Distribution Matching (RT-DM) paradigm to ”softly” integrate reward signals into the distillation process.
+Specifically, we utilize the reward score as a weighting factor for the DMD loss.
+This formulation enables a more flexible optimization objective compared to direct reward maximization.
+This is because the reward score serves as a static weighting factor, we avoid the necessity of backpropagating through the multi-step denoising process, which allows us to evaluate the reward on the complete trajectory of the few-step model (e.g., 4 steps) without the computational overhead of a large gradient graph, bypassing the need for single-step $\bm{x}_{0}$ approximations to make the reward score more precise and highly valuable as a reference for optimization efforts.
+More profoundly, we are changing which distribution the student learns by smoothly integrating reward signals to make the student not learn toward the vanilla $P_{\text{real}}$, but toward a reward-tilted teacher distribution defined proportionally to $r(x_{0})p_{\text{real}}$.
+The gradient is formulated as:
+$\nabla_{\!\theta}\,\mathcal{L}_{\text{RT-DM}}$ $=\mathbb{E}_{t}\!\left[\nabla_{\!\theta}\operatorname{KL}(p_{\text{fake},t}\|(p_{\text{real},t})r(\bm{x}_{0}))\right]$ (4)
+$=-\mathbb{E}_{t}\!\left[\int\!e^{\frac{r(\bm{x}_{0})}{\left\|r(\bm{x}_{0})\right\|}}\Bigl(s_{\text{real}}\bigl(F_{t})\bigr)-s_{\text{fake}}\bigl(F_{t}\bigr)\Bigr)\frac{\mathrm{d}G_{\theta}(z)}{\mathrm{d}\theta}\,\mathrm{d}z\right]\!.$
+To evaluate the effectiveness of the proposed soft integration, we compare three configurations in Fig.
+4: (i) DMD Only (Blue); (ii) RT-DM for the entire duration (Red); and (iii) RT-DM followed by joint DMD+RL (Green).
+As shown in Fig.
+4(a) and (b), the RT-DM paradigm (Red/Green) provides a significant ”warm start” effect, accelerating preference alignment in the early phase (Stage 1) compared to the DMD-only baseline.
+This confirms that weighting the DMD loss with reward scores successfully incorporates preference signals even when few-step generations are still maturing, avoiding the noise instability of direct RL.
+However, the performance of the ”RT-DM only” approach (Red) eventually saturates.
+We hypothesize that since RT-DM primarily acts by re-weighting the teacher’s distribution gradients, its ability to drive the model toward high-reward regions is inherently constrained by the teacher’s manifold.
+To further unlock the model’s potential, we introduce a more direct reward maximization phase (Green) in Stage 2.
+By transitioning to joint DMD+RL optimization once the model has reached a stable baseline, we can leverage direct RL gradients to aggressively push the performance beyond the initial plateau.
+As a result, the joint approach achieves the highest test rewards and DPG scores, demonstrating that RT-DM serves as an ideal bridge to transition from pure distillation to direct preference alignment.
+2.5 Annealed Distributional Overlap Maximization for Better Initial Distillation
+During the initial phase of distillation, DMD encounters a classic ”cold start” challenge: the synthetic samples generated by the few-step student exhibit substantial divergence from the real-world pre-training samples.
+This profound disjointness between $P_{\text{fake}}$ and $P_{\text{real}}$ impedes the teacher model’s ability to estimate reliable scores, rendering the gradient $\nabla\log P_{\text{real}}-\nabla\log P_{\text{fake}}$ unreliable for optimization.
+To mitigate this, we introduce a unified principle of annealed distributional overlap maximization.
+We propose two complementary techniques to artificially increase the overlap between $P_{\text{fake}}$ and $P_{\text{real}}$ early in training, progressively relaxing them as the model converges.
 Dynamic Distribution Guidance (DynaDG).
-In the initial stage of training, due to the significant disparity in image quality (the generator is unable to generate qualified images), the distribution of these generated samples ($P_{fake}$) is completely different from that of the training samples of the real score estimator ($P_{real}$), which finally leads to the inaccuracies in the KL divergence approximation as we show in Figure 6 (left).
-This substantial divergence comes from sparse or non-existent overlap between $P_{real}$ and $P_{fake}$, rendering the gradient $\nabla\log P_{real}-\nabla\log P_{fake}$ unreliable for effective optimization.
-Inspired by a recent study [magic141], we introduce a LoRA set [lora] and inject it not only into the fake score estimator to learn the rapidly evolving distribution of the few-step generator but also strategically into the real score estimator.
-The core idea behind our DynaDG is to dynamically adjust the distribution estimated in the real score branch.
-By injecting LoRA into the real score estimator with a smaller scale, we can effectively “pull” the perceived $P_{real}$ towards the nascent $P_{fake}$ distribution, creating a region of “more overlap” as depicted in Figure 6 (right).
-This increased overlap ensures that the estimated score difference, $\nabla\log P_{real}-\nabla\log P_{fake}$, consistently provides a “right direction estimation and optimization” signal, even when the distributions are initially far apart.
-Crucially, this LoRA scale for the real score estimator gradually weakens as training progresses.
-This mechanism allows the real score estimator to progressively revert to accurately representing the true $P_{real}$ as the generator improves and $P_{fake}$ naturally converges towards $P_{real}$.
-Figure 6: Illustration for our Dynamic Distribution Guidance.
+DynaDG maximizes overlap primarily by actively shifting the target distribution.
+The poor initial Image quality causes $P_{\text{fake}}$ to be disjoint from $P_{\text{real}}$ (Figure 6, left).
+To solve this, we inject trainable LoRA modules [lora] into the real score estimator to ”pull” the perceived $P_{\text{real}}$ toward the nascent $P_{\text{fake}}$ (Figure 6, right).
+This ensures a reliable optimization signal by bridging the manifold gap.
+As training progresses, we anneal the LoRA scale in the teacher model toward zero, allowing the real score estimator to seamlessly revert to the true $P_{\text{real}}$ once $P_{\text{fake}}$ is safely anchored.
+Figure 5: Impact of dynamic training strategies.
+(a) Qualitative comparisons show that our dynamic strategies facilitate faster global structure building (highlighted in red boxes).
+(b) Quantitative DPG scores further confirm that DynaDG and DynaRS provide more reliable optimization signals, consistently outperforming the baseline throughout Stage 1.
+Better to zoom in to check the effect.
 Dynamic Renoise Sampling (DynaRS).
-Another important operation in DMD is renoising, which renoises the samples generated by the generator for estimating the score (i.e., the distribution) associated with the specific noise level.
-The core idea of our DynaRS is to initially bias the sampling of renoise levels $t$ in $F_{t}$ towards higher values, and then gradually transition to a uniform sampling like DMD dose.
-This has two advantages for a faster and better cold start.
-First, when biasing to high noise, the forward operator makes noise largely occupy the image, so even bad samples produced by the generator in the initial stage lie in a similar regime for both real and fake estimators, thus improving score reliability and overlap.
-Second, the estimated score at a higher noise level is more about global structure [lsast, prospect], which can guide the generator to first capture global structure.
-As the capacity of the generator improves, the renoise bias is progressively reduced, allowing the model to learn finer details by sampling from a wider range of noise levels.
-3.4 Distinction from Prior Few-step RL Works
-Some prior works also try to introduce RL for the few-step diffusion model [hypersd, ladd, hyperbagel, hypernoise], but they mainly perform it after the distillation, and the problem mentioned by Hyper-SD [hypersd] of quickly deviating from the original domain still occurs.
-In this work, we illustrate the rationale for conducting the two simultaneously and how DMD loss can effectively regularize RL to mitigate the above problems.
-Meanwhile, although pioneer work like Diff-Instruct ++ [diffinstruct++] also introduces RL in the distillation process, our motivation and analysis are distinct, with more intuitive illustrations of the mutual benefits of the two, the verification of different diffusion models and RL algorithms at a large scale, as well as the dynamic cold start stage before incorporating RL.
+While DynaDG moves the distributions, DynaRS maximizes overlap by exploiting regions where distributions naturally intersect.
+By heavily biasing the sampling of renoise levels $t$ toward higher noise values at the start of training, we ensure that both initial samples and reference samples are dominated by Gaussian noise, placing them in a similar regime.
+At these high noise levels, the score estimator provides highly reliable gradients that emphasize global structure [lsast, prospect].
+As the student generator improves, the renoise bias is progressively annealed to uniform sampling, allowing the model to capture fine-grained textural details from a wider range of noise levels.
+Figure 6: Illustration for Dynamic Distribution Guidance.
+As shown in Figure 5 (b), our dynamic training strategies significantly enhance the initial training phase.
+Qualitative results presented in Figure 5 left further corroborate these findings, showing a faster global structure building when DynaDG and DynaRS are employed.
 3 Experiments
 3.1 Experimental Setup
-We provide a brief overview of our training and evaluation schemes here.
-For training, we adopt denoising-based model (SDXL-Base [sdxl]) and flow-based models (SD3-Medium [sd3], SD3.5-Large [sd35]) for distillation, using prompts from t2i-2M [t2i-2M].
+For training, we adopt denoising-based model (SDXL-Base [sdxl]) and flow-based models (SD3-Medium [sd3], SD3.5-Large [sd35]) for distillation, using prompts from t2i-2M.
 Additionally, we use ReFL [refl] with DFN-CLIP [dfnclip] and HPSv2.1 [hpsv2] as reward models by default.
@@
 Next, in order to obtain a more comprehensive evaluation, we also compare our distilled models with their teachers on two popular benchmarks, DPG_Bench [dpg] and GenEval [geneval].
-Finally, we conduct both qualitative and quantitative ablation experiments to further verify our method’s effectiveness.
-Table 1: System-level comparison against state-of-the-art methods.
-denotes our reproduced results.
-Best performance is marked in Bold.
-Image-Free denotes whether the training requires external image data or data generated by the multi-step teacher.
+Noted that there are some differences from the settings we made in the Method section.
+We train 1.5K steps in stage1 and stage 2, respectively, to achieve better performance.
+Table 2: System-level comparison against state-of-the-art methods. denotes our reproduced results based on the official code. Best performance is marked in Bold.
 […]
 SDXL-Base
-Base-Model (CFG = 7.0) 25 50 34.7588 5.6480 22.1085 27.1477 -
-LCM [lcm] (CFG = 8.0) 1 2 28.4664 5.1026 20.0603 17.6837 $\times$
-Lightning [sdxl-lightning] 1 1 32.0283 5.6761 21.4868 26.3615 $\times$
-DMD2 [dmd2] 1 1 34.3046 5.6238 21.7293 26.5986 $\times$
-DMDR (ours) 1 1 35.4835 6.0483 22.5424 31.1442 $\checkmark$
-DMD2 [dmd2] 4 4 34.5169 5.7043 22.1546 28.5655 $\times$
-DMDR (ours) 4 4 35.2940 5.9857 22.6268 32.8678 $\checkmark$
+Base-Model 50 - 34.7588 5.6480 22.1085 27.1477
+ReFL [refl] 50 RL 35.2107 5.9045 22.8076 33.1874
+LCM [lcm] 2 Distill 28.4664 5.1026 20.0603 17.6837
+Lightning [sdxl-lightning] 1 Distill 32.0283 5.6761 21.4868 26.3615
+DMD2 [dmd2] 1 Distill 34.3046 5.6238 21.7293 26.5986
+DMDR (ours) 1 Distill & RL 35.6241 6.1184 22.8498 32.0065
+DMD2 [dmd2] 4 Distill 34.5169 5.7043 22.1546 28.5655
+DMD2-PSO [pso] 4 Distill then RL 34.0128 5.8032 22.2644 29.8732
+DMDR (ours) 4 Distill & RL 35.4940 6.0324 22.7122 32.9832
 SD3-Medium
-Base-Model (CFG = 7.0) 25 50 34.9025 5.5942 22.1801 28.4021 -
-Hyper-SD [hypersd] (CFG = 5.0) 4 8 32.0234 5.2489 20.2831 22.4544 $\times$
-DMD2 [dmd2] 4 4 33.9421 5.6137 21.7688 27.3675 $\times$
-Flash [flashdiff] 4 4 34.2634 5.6702 21.5921 26.6542 $\times$
-TDM [tdm] 4 4 34.0301 5.6250 22.0010 27.7522 $\checkmark$
-DMDR (ours) 4 4 34.9542 5.8462 22.3578 31.8979 $\checkmark$
+Base-Model 50 - 34.9025 5.5942 22.1801 28.4021
+ReFL [refl] 50 RL 35.1851 5.7821 22.0064 32.0745
+DMD2 [dmd2] 4 Distill 33.9421 5.6137 21.7688 27.3675
+Flash [flashdiff] 4 Distill 34.2634 5.6702 21.5921 26.6542
+TDM [tdm] 4 Distill 34.0301 5.6250 22.0010 27.7522
+Hyper-SD [hypersd] 8 Distill then RL 32.0234 5.2489 20.2831 22.4544
+DMDR (ours) 4 Distill & RL 35.0142 5.8876 22.4892 32.1145
 SD3.5-Large
-Base-Model (CFG = 3.5) 25 50 35.5509 5.7014 22.4856 28.8135 -
-LADD [ladd] 4 4 35.0480 5.4514 22.2451 27.8470 $\times$
-DMDR (ours) 4 4 35.8647 6.0284 22.8859 32.4724 $\checkmark$
-Table 2: Quantitative comparison on DPG_Bench [DPG] of our four-step model against its multi-step teacher.
-Model Overall Global Entity Attribute Relation Other
-SDXL-Base
-Teacher 74.65 83.27 82.43 80.91 86.76 80.41
-Ours 76.44 83.70 82.54 83.66 84.75 83.49
-SD3-Medium
-Teacher 84.08 87.90 91.01 88.83 80.70 88.68
-Ours 84.96 90.46 90.05 90.52 87.12 91.23
-SD3.5-Large
-Teacher 84.12 91.48 90.22 87.81 91.20 89.49
-Ours 85.30 90.46 90.50 90.66 87.40 90.18
+Base-Model 50 - 35.5509 5.7014 22.4856 28.8135
+LADD [ladd] 4 Distill then RL 35.0480 5.4514 22.2451 27.8470
+DMDR (ours) 4 Distill & RL 35.9757 6.1541 22.9072 32.7368
+Method for comparison.
+We compare our method against several categories of existing work, including foundational multi-step base models [sdxl, sd3, sd35], RL only approaches [refl], distillation only methods [lcm, sdxl-lightning, dmd2, flashdiff, tdm], distillation then RL frameworks [hypersd, pso].
+We do not compare with methods like Flash-DMD [flash-dmd] and Diff-Instruct++ [diffinstruct++] that do not provide model weights or open source training code because we cannot guarantee a fair and reproducible evaluation.
 3.2 Main Results
@@
 Notably, our method demonstrates effectiveness across various architectures (e.g., UNet, Transformer), model sizes (e.g., 2B, 8B), and paradigms (e.g., flow-based, denoising-based), highlighting the broad applicability and universality of the proposed approach.
-Additionally, we provide qualitative comparisons in Figure 5 and Appendix B, where our method produces images characterized by superior quality and enhanced prompt coherence.
-To further verify our method’s effectiveness and provide a more comprehensive comparison between the few-step distillation model and the multi-step teacher model, we conduct the comparison on two commonly used benchmarks [DPG, geneval], Table 2 shows our four-step model consistently outperforms its multi-step teacher on DPG_Bench across all base models, lifting the overall score from 74.65 to 76.44 for SDXL-Base, 84.08 to 84.96 for SD3-Medium, and 84.12 to 85.30 for SD3.5-Large.
-Meanwhile, on GenEval (Table 3) the same trend holds: overall scores improve for SDXL (0.55 → 0.56), SD3 (0.62 → 0.64), and SD3.5 (0.71 → 0.72).
-These results indicate that DMDR has successfully freed the student model from the constraints of the multi-step teacher model and stimulated its capabilities during the distillation process.
-Table 3: Quantitative comparison on GenEval [geneval] of our four-step model against its multi-step teacher.
+Additionally, we provide qualitative comparisons in Figure 7, where our method produces images characterized by superior quality and enhanced prompt coherence.
+Table 3: GenEval Comparison: 4-step (ours) $vs.$ its multi-step teacher.
 Model Overall Single Two Count. Colors Pos. Attr.
@@
 Teacher 0.55 0.98 0.74 0.39 0.85 0.15 0.23
-Ours 0.56 0.99 0.76 0.42 0.84 0.11 0.24
+Ours 0.57 0.99 0.76 0.44 0.85 0.12 0.25
 SD3-Medium
 Teacher 0.62 0.98 0.74 0.63 0.67 0.34 0.36
-Ours 0.64 0.99 0.84 0.54 0.81 0.25 0.44
+Ours 0.65 0.99 0.84 0.57 0.81 0.27 0.45
 SD3.5-Large
 Teacher 0.71 0.98 0.89 0.73 0.83 0.34 0.47
-Ours 0.72 0.99 0.93 0.68 0.80 0.30 0.59
-Figure 7: Ablation examples of two dynamic training strategies. Better to zoom in to check the effect.
+Ours 0.72 0.99 0.93 0.69 0.80 0.31 0.59
+Table 4: DPG_Bench Comparison: 4-step (ours) $vs.$ its multi-step teacher.
+Model Overall Global Entity Attribute Relation Other
+SDXL-Base
+Teacher 74.65 83.27 82.43 80.91 86.76 80.41
+Ours 76.48 83.72 82.58 83.69 84.76 83.54
+SD3-Medium
+Teacher 84.08 87.90 91.01 88.83 80.70 88.68
+Ours 85.30 90.49 90.07 90.56 87.17 91.23
+SD3.5-Large
+Teacher 84.12 91.48 90.22 87.81 91.20 89.49
+Ours 85.33 90.47 90.53 90.68 87.44 90.20
+Figure 7: Visual comparison between the teachers, selected competing methods [dmd2, hypersd, ladd], and ours.
+All images are generated using identical noise.
+Our model produces images with superior quality and prompt coherence.
+To further verify our method’s effectiveness and provide a more comprehensive comparison between the few-step distillation model and the multi-step teacher model, we conduct the comparison on two commonly used benchmarks [dpg, geneval], Although we do not perform RL optimization on the metric of these two benchmarks, table 4 and  4 show that the overall score of our four-step model consistently outperforms its multi-step teacher on DPG_Bench and GenEvalacross all base models.
+These results indicate that DMDR has successfully freed the student model from the constraints of the multi-step teacher model and stimulated its capabilities during the distillation process.
+Table 5: Ablation on different components. Default settings are marked in purple.
+(a) Ablation on different reward models.
+Method CLIP Aesthetic Pick HP DPG
+Score Score Score Score overall
+Pick [pickscore] 34.3403 5.9021 22.9674 31.7665 84.86
+AE [laion5b] 34.1748 6.1024 22.3858 31.6620 84.61
+CLIP [clipscore] 35.2087 5.5832 22.0644 28.0507 85.10
+HPS [hpsv2] 34.4088 5.9021 22.5108 32.9065 85.04
+CLIP + HPS 35.0142 5.8876 22.4892 32.1145 85.30
+(b) Ablation on RL algorithm in stage 2.
+Method CLIP Aesthetic Pick HP DPG
+Score Score Score Score overall
+stage 1 init 33.3648 5.5744 21.0070 28.8371 82.4371
+w/ RL (ReFL) 35.0142 5.8876 22.4892 32.1145 85.30
+w/ RL (DPO) 34.0234 5.8340 21.9844 30.8352 85.00
+w/ RL (GRPO) 34.6675 5.9324 22.3248 31.4474 85.08
+(c) Ablation on Value of $\lambda_{\text{rl}}$.
+Value CLIP Aesthetic Pick HP DPG
+Score Score Score Score overall
+0.1 34.7422 5.8508 22.3064 31.6565 85.24
+0.5 35.0142 5.8876 22.4892 32.1145 85.30
+1.0 35.0087 5.84430 22.4709 32.8982 85.04
+2.0 35.0322 5.7806 22.1098 32.7683 82.48
 3.3 Ablation Study
-We investigate our method’s properties below. We use SD3-Medium with $512\times 512$ resolution by default for a fast ablation.
-Effect of Dynamic Cold Start Stage.
-As shown in Table 4.3, our dynamic cold start strategies significantly enhance the initial training phase.
-Firstly, introducing DynaDG leads to a notable improvement.
-By dynamically adjusting the $P_{real}$ to create more overlap with the nascent $P_{fake}$ distribution, DynaDG provides more reliable gradient signals.
-Quantitatively, this translates to a CLIP Score increase from 31.56 to 32.79 (+1.23) and an HP Score increase from 27.01 to 27.57 (+0.56) at 1000 iterations.
-Similar gains are observed at 2000 iterations.
-Further incorporating DynaRS, which initially biases renoise sampling towards higher noise levels, guiding the generator to learn global structures before progressively shifting to finer details, further boosts performance.
-To highlight the importance of the dynamic aspect of our strategies, we evaluate a variant denoted as -Dynamic.
-This variant removes the dynamic scaling of LoRA in DynaDG (e.g., using a fixed LoRA scale) and employs biased renoise sampling in the whole training process.
-While it still outperforms vanilla_DMD, it consistently underperforms the full dynamic approaches.
-This clearly indicates that the adaptive, progressive nature of our DynaDG and DynaRS mechanisms is crucial for maximizing performance during the cold start phase.
-Qualitative results presented in Figure 7 further corroborate these findings, showing a faster global structure building when our dynamic cold start strategies are employed.
-Table 4: Ablation on our dynamic cold start stage.
-Method 1000 iter 2000 iter
-CLIP Score HP Score CLIP Score HP Score
-vanilla_DMD 31.5608 27.0051 32.5492 28.4541
-   +DynaDG 32.7905 27.5723 33.5045 29.0363
-   +DynaRS 33.1657 27.8276 33.6432 29.1157
-   -Dynamic 32.6562 27.2641 33.0245 28.8459
-Effect of DMDR with Different RL Algorithms.
-We choose three representative RL methods [refl, diffusiondpo, flowgrpo] to illustrate the efficacy of combining distillation with various RL techniques in DMDR, the results are provided in Table 4.3.
-Firstly, we observe that the initial model state after our dynamic cold start stage already provides a good baseline performance.
-Training with “w/ only Distill.”, as we illustrated above, this would be inherently capped by the teacher’s distribution, resulting in a very minor improvement over the initial state.
-Next, when considering introducing different RL algorithms (the regularization originally used in these methods has also been added), although these show improvements in some metrics, they do not consistently surpass the “only Distill” baseline across all scores.
-Moreover, the reward hacking occurs (we will analyze it specifically in the next part).
-Crucially, the rows corresponding to “w/ Distill.
-+ RL” consistently demonstrates superior performance across all evaluated metrics against the counterparts.
-Figure 8 also shows superior visualization results when DMD and RL are conducted together.
-These validate our core insight that RL unlocks the performance potential of DMD, allowing the student model to surpass the teacher’s inherent limitations by optimizing for external reward signals.
-Simultaneously, DMD effectively regularizes the RL process, preventing reward hacking and ensuring that the student’s distribution remains robust and comprehensive, as evidenced by the consistent improvements across diverse metrics with different RL algorithms.
-Table 5: Ablation on RL and DMD. We init the model after cold start for 2K iteration then continue training in different ways.
-Method CLIP Aesthetic Pick HP
-Score Score Score Score
-init 33.6432 5.6124 21.0489 29.1157
-w/ only Distill. 33.6738 5.6248 21.6376 29.1389
-w/ only RL (ReFL) 33.1897 5.8841 22.3008 31.2714
-w/ Distill. + RL (ReFL) 34.6249 6.1813 22.7578 32.8979
-w/ only RL (DPO) 33.9554 5.7893 21.7643 29.5819
-w/ Distill. + RL (DPO) 33.9632 5.9710 21.9865 30.5994
-w/ only RL (GRPO) 33.9095 5.7876 21.7956 29.4429
-w/ Distill. + RL (GRPO) 34.0055 5.8256 22.0120 30.6248
-Figure 8: Ablation examples of different RL algorithms with DMD. Better to zoom in to check the effect.
-Furthermore, an interesting observation is that using the ReFL style RL method achieves overall better performance, and our analysis is that: traditional ReFL for multi-step models ignores earlier sampling steps and only trains the last few steps before the output image [refl] because the large computational graph generated by the gradient of the reward model and the multi-step denoising process would result in a huge memory consumption.
-Subsequent works [srpo, deepreward] mainly focus on saving forward noise trajectory and sampling a subset of steps to optimize to remedy this dilemma, although this can allow the gradient to return to the early denoising step, it is also prone to error accumulation.
+In this section, we conduct a systematic investigation into the key components of DMDR. Unless otherwise specified, all experiments use the SD3-M base model with a 4-step sampling configuration.
+Versatility across reward models.
+A key advantage of ourusing a DINOv2-based reward model to maximize classification accuracyDMDR is the ability to steer the distillation process toward diverse human preferences.
+As shown in Table 5(a), we evaluate the framework using different reward models (RMs) targeting specific attributes.
+The results provide a clear insight: optimizing with a specific RM leads to a significant performance gain in its corresponding metric, confirming that DMDR effectively injects the desired preference signal into the few-step manifold.
+Furthermore, by integrating multiple rewards (CLIP + HPS), the model achieves a more balanced and superior generative performance, similar to the findings in DanceGRPO [dancegrpo].
+This flexibility allows users to customize the distillation ”flavor” according to specific downstream tasks.
+Compatibility with RL paradigms.
+To verify the algorithmic universality of our unified framework, we compare three distinct RL strategies in the Stage 2 joint optimization phase: ReFL [refl], DPO [diffusiondpo, pso], and GRPO [dancegrpo, flowgrpo].
+As shown in Table 5(b), all three algorithms yield substantial improvements over the ”Stage 1 init” baseline across all evaluation dimensions.
+Furthermore, an interesting observation is that using the ReFL in DMDR achieves overall better performance, and our analysis is that: traditional ReFL for multi-step models ignores earlier sampling steps and only trains the last few steps before the output image [refl] because the large computational graph generated by the multi-step denoising process would result in a huge memory consumption.
+Subsequent solutions [srpo, deepreward] mainly focus on saving forward noise trajectory and sampling a subset of steps to optimize, although this can allow the gradient to return to the early denoising step, it is also prone to error accumulation.
 On the contrary, in our few-step model, the model is trained to predict a clean image at each step’s state [dmd2], which enables the gradient of ReFL to directly return to the initial state, thus obtaining an effective optimization.
-Figure 9: Reward curves comparison during the RL stage. Better to zoom in to check the difference.
-Effect of DMDR to Mitigate Reward Hacking.
-We follow RewardDance [rewarddance] to present the reward curves during training illustrated in Figure 9.
-It is evident that employing DMD loss as a form of regularization (right) demonstrates significantly greater variance throughout the training process compared to the traditional approach (left).
-This increased diversity in reward signals suggests that models trained using our method exhibit enhanced robustness against reward hacking.
-This assertion is further supported by Figure 8, which shows that in the absence of DMD loss, utilizing ReFL results in a rasterized appearance of images, indicating the occurrence of reward hacking.
-5 Limitation and Future Work
-We discuss some limitations we realized during the development of DMDR. We hope these will serve as a catalyst for innovative ideas and advancements in future research.
+The balancing Role of DMD Regularization.
+The synergy between distillation and RL is governed by the balancing coefficient $\lambda_{\text{rl}}$ in our joint objective.
+Table 5(c) illustrates the trade-off between reward maximization and distributional stability.
+When $\lambda_{\text{rl}}$ is set too low; the model remains constrained by the teacher’s manifold, limiting the potential for preference alignment.
+Conversely, as it increases beyond an optimal threshold, we observe a noticeable decline in a general capacity (reflected by the DPG overall score), despite high reward scores.
+This trend provides a crucial insight: the DMD loss serves as a vital structural regularizer.
+It prevents the model from ”reward hacking”(We also showed in Figure 3.
+The optimal balance ensures that the model pushes the boundaries of quality while staying anchored to the teacher’s robust generative priors.
+4 Limitation and Discussion
 Quality vs.
 Diversity.
-One notable thing is the trade-off between quality and diversity.
-Although models trained from our method show superior quality and prompt coherence, the diversity is reduced like shown in Table 6.
-We believe that this is a common feature of score distillation like DMD and RL, that is, sacrificing uncertainty in sampling in exchange for stable and desired results, not a limitation unique to our approach.
-The reduced diversity caused by DMD, we believe, will also have some impact on some online RL methods that require a group of samples for ranking then optimization.
-This is also reflected in the slightly poorer improvement of GRPO and DPO compared to ReFL in the Table 4.3.
-Further investigation into finding the optimal balance between quality and diversity is left for future work.
-Table 6: Diversity evaluation on ShareGPT-4o-Image [sharegpt4o-image].
-DMD DMD2 Ours Teacher
-[…]
-Scaling of Reward Model.
-Another notable thing is the scaling of the reward model in DMDR.
-Our experimental results demonstrate that our method enables the few-step model to be able to surpass its teacher model.
-However, as indicated in the Table 2 and 3, certain metrics exhibit a decline after the distillation.
-We attribute this degradation primarily to inherent limitations and potential biases present in the open-source reward model we employed.
-Notably, RewardDance [rewarddance] has shown that scaling up the reward model yields consistent improvements for RL in multi-step diffusion models.
-We posit that similar benefits would extend to few-step models with our method, but require more computation resources, which we leave for future work.
-6 Conclusion
-This work introduces DMDR, which combines Distribution Matching Distillation (DMD) with reinforcement learning (RL) in one framework.
-We show that DMD and RL can be trained simultaneously, with RL enabling the student model to surpass the teacher and DMD loss regularizing RL to prevent reward hacking.
-We also propose dynamic training strategies (DynaDG and DynaRS) to enhance the initial distillation process.
-Experiments demonstrate state-of-the-art performance among few-step methods and overall superior performance against multi-step teachers.
-Appendix
-Appendix A More Related Works
+To further analyze the behavior of DMDR, we conduct class-conditional experiments on ImageNet $\times256$ using the SiT [sit] backbone with representation alignment pretraining ckpt [repa, sra] as initialization for 1-step generation.
+As shown in Table 6, a fundamental trade-off between generative quality and diversity emerges.
+Our results reveal that both Classifier-Free Guidance (CFG) (this also observed in Decoupled-DMD [decoupled-dmd]) and Reinforcement Learning (here we use a DINOv2-based [dinov2] reward model to maximize classification accuracy) act as ”density sharpeners” during the distillation process.
+This phenomenon suggests that DMDR effectively ”sacrifices” the teacher’s distributional uncertainty in exchange for the stability and quality required for few-step sampling.
+We believe this reduction in diversity is not a limitation unique to our approach, but a characteristic of score-based distillation and preference alignment, which steer the model toward high-reward, high-confidence regions.
+This trade-off is often desirable in practical few-step applications where sample excellence and prompt adherence are prioritized over exhaustive distribution coverage.
+But at the same time, further investigation into finding the optimal balance between quality and diversity is left for future work.
+Table 6: 1-step SiT DMDR Results on ImageNet $256\times 256$.
+We report the performance with different CFG scales and RL optimization.
+Metrics include FID [fid], Recall [pre-rec], Inception Score (IS) [is], and Precision [pre-rec].
+CFG RL FID $\downarrow$ Recall $\uparrow$ Inception Score $\uparrow$ Precision $\uparrow$
+No No 2.13 0.64 232.15 0.77
+1.5 No 5.20 0.48 387.24 0.88
+4 No 15.70 0.11 453.65 0.86
+No Yes 6.95 0.40 416.01 0.90
+1.5 Yes 13.43 0.24 494.43 0.93
+5 Related Work
+Distribution matching distillation.
+Distribution Matching Distillation [dmd] (DMD) is the first work to successfully apply the principle score-based distillation [score-base-diff, prolificdreamer, dreamfusion] to large-scale text-to-image models.
+Intuitively, it strives to ensure that any sample realized by the student at a given noise level occurs with exactly the same probability as it would under the teacher’s distribution, thereby preserving the multi-step generative priors in the few-step model.
+From then on, a lot of follow-up work emerged [dmd2, dmdx, tdm, sd3.5flash, f-distill, flash-dmd, twinflow], for example, DMD2 [dmd2] employs a discriminator to align the student model distribution with a specific target distribution like GAN [gan] does; TDM [tdm] incorporates DMD loss in the sampling process of the student model for better alignment; ff–distill [f-distill] replaces the original reverse Kullback–Leibler (KL) divergence to the proposed $f$-divergence for covering different divergences with different properties.
+Our work also starts with DMD, but we don’t focus on how to better “imitate” the teacher.
+Instead, we aim to incorporate reinforcement learning in the distillation process to enable a more controllable and preference-aligned distillation.
+Reinforce learning for diffusion models.
+Reinforce learning helps to align diffusion models to human preferences by training a reward model and using it to guide generation [ddpo, flowgrpo, refl, srpo, diffusiondpo, dancegrpo, diffusionnft].
+There are many algorithms to conduct RL, for example, DDPO [ddpo] adapts PPO via image log-likelihoods; ReFL [refl] bypasses likelihoods by optimizing outputs with frozen-reward gradients; Diffusion-DPO [diffusiondpo] adapts DPO to diffusion for paired human preference data; and recent GRPO extensions [flowgrpo, dancegrpo] use GRPO to diffusion models by coupling the training loss with SDE samplers.
+Although significant progress has been made in RL for diffusion models, most of the work has focused on multi-step models.
+Here, we find that these algorithms are also adapted to the few-step model when carried out in conjunction with distribution matching distillation.
 Trajectory-based distillation.
@@
 TDM [tdm] and rCM [rcm] combine DMD [dmd] and LCM [lcm] together to remedy the quality issues of LCM.
-Appendix B Additional Visual Comparison Results
-In this section, we provide additional visual comparisons for our 4-step distilled models against teacher models and other methods. The results are shown in Figure 10, Figure 11:
-Figure 10: Additional visual comparison in SDXL-Base. All images are generated using identical noise and text prompts. Our model produces images with superior quality and prompt coherence.
-Figure 11: Additional visual comparison in SD3-Medium. All images are generated using identical noise and text prompts. Our model produces images with superior quality and prompt coherence.
-Figure 12: Additional visual comparison in SD3.5-Large. All images are generated using identical noise and text prompts. Our model produces images with superior quality and prompt coherence.
+Reward models for diffusion post-train.
+Unlike Large Language Models (LLMs), where reinforcement learning (RL) can often leverage verifiable, rule-based rewards [deepseekmath, deepseekr1, glm] (e.g., code execution success or mathematical correctness), diffusion-based image generation is inherently subjective.
+The evaluation of generated images relies heavily on multifaceted criteria such as aesthetic quality, instruction following, and so on, which are difficult to quantify through rigid, rule-based functions.
+Consequently, the development of robust reward models (RMs) has become a critical prerequisite for effective RL-based alignment in diffusion models [refl, rewarddance, clipscore, hps, visionreward, q-align].
+Early methods, such as CLIP-score [clipscore], primarily focus on measuring the semantic alignment between the text prompt and the generated image.
+To address the need for visual fidelity and artistic appeal, metrics like Aesthetic Score [laion5b] and HPS (Human Preference Score) [hps, hpsv2, hpsv3] have been introduced to capture the nuance of human aesthetic judgment.
+More recently, the reward models become unified [unified-reward, visionreward, unified-p-reward] and aggregate multi-dimensional preferences—ranging from low-level texture details to high-level semantic instruction following—into a unified scoring function.
+While our primary contribution lies in proposing a novel algorithm that integrates few-step distillation with reinforcement learning, our empirical results highlight the pivotal role of the reward model in our pipeline.
+6 Conclusion
+We introduces DMDR, a unified framework that transforms Distribution Matching Distillation (DMD) and Reinforcement Learning (RL) from independent, sequential stages into a synergistic training pipeline.
+We demonstrate that joint optimization is mutually beneficial: RL steers distillation toward high-reward regions to break the ”teacher ceiling,” while DMD regularizes RL to effectively mitigate reward hacking.
+By employing a two-stage strategy—incorporating RT-DM and dynamic training followed by joint optimization—DMDR achieves state-of-the-art few-step performance across diverse architectures.
+Ultimately, our approach enables few-step models to not only match but consistently surpass the visual quality and prompt adherence of their original multi-step teacher models.